OpenStudy (anonymous):

Tom has a collection of 30 CDs and Nita has a collection of 18 CDs. Tom is adding 1 CD a month to his collection while Nita is adding 5 CDs a month to her collection. Find the number of months after which they will have the same number of CDs.

OpenStudy (anonymous):

T = 30 + m N= 18 + 5m -> T = N ... = ....

OpenStudy (anonymous):

x = 3.

OpenStudy (anonymous):

30 + x = 18 + 5x. Solve for x to get x = 3.

OpenStudy (anonymous):

It will take 3 months in order for Nita will have the same number of CDs as Tom does!